# Deep Learning for VWAP Execution in Crypto Markets:  Beyond the Volume Curve

### v1

@@ 6.3 Experimental Setup
The strategies are evaluated using the dataset described in Section 2. The dataset consists of hourly data from five major cryptocurrencies (BTC, ETH, BNB, ADA, and XRP) traded on Binance perpetual contracts from inception until 1 July 2024. The data is partitioned chronologically, with the last 20% reserved as the test set to ensure a true out-of-sample evaluation. A validation set, comprising 20% of the remaining data, is randomly selected from the first 80% of the period to guide model selection and prevent overfitting to any specific market regime. For each cryptocurrency and strategy, 30 training runs are performed to compute both the mean and standard deviation of the performance metrics. In these experiments, a lookback period of 120 time steps is used to predict allocations for 12 periods ahead. Additional experiments with modified parameters—one with a 6-step prediction horizon and a shorter lookback period, and another with a 48-step prediction horizon—are provided in the Appendix. The results from these experiments confirm that the framework is effective across different timeframes and that the underlying concepts generalize well. The input features provided to the models include volumes (scaled by dividing by the two-week rolling average and shifted by the sum of the lookback and lookahead periods to avoid forward-looking bias), the hour of the day, the day of the week, and returns computed on the VWAP price of each bin (with a value of 0 assigned when no volume is traded). As a target, the volumes in the lookahead period are scaled by dividing by the sum of the volumes in the lookback period. This scaling produces a more stable target series and does not affect the VWAP loss training by construction, although it does impact the calibration of standard models, particularly the linear regression model.
@@ 6.5 Empirical Results
Table 2 summarizes the performance of various VWAP execution strategies for a 12-step-ahead allocation with a 120-step lookback window. The table reports the mean and standard deviation of the Absolute VWAP Loss (scaled by $10^{-2}$), Quadratic VWAP Loss (scaled by $10^{-4}$), the $R^{2}$ score for the volume curve, and the training time (in seconds) for each strategy, across five assets (BTC, ETH, BNB, ADA, XRP). For Bitcoin (BTC), the naive strategy yields an absolute VWAP loss of 0.1587 and a quadratic VWAP loss of 0.0878. In contrast, the StaticVWAP model optimized using the absolute loss reduces these metrics to 0.1197 and 0.0502, respectively—indicating improvements of roughly 25% in absolute loss and 43% in quadratic loss. When optimizing for the quadratic loss, the StaticVWAP model further improves the quadratic loss to 0.0474, although the corresponding $R^{2}$ value becomes more negative (–0.3616), suggesting that an optimal execution strategy does not necessarily mimic the raw volume curve. The fixed volume curve approach, while yielding a relatively better $R^{2}$ (–0.4670 for quadratic loss), comes at the cost of a higher absolute loss (0.1275) and a training time that is more variable across assets.
@@
Dynamic Linear BTC N/A 0.134587 0.000000 0.076583 0.000000 0.178439 0.000000 0.384068 0.000000
Static Linear BTC N/A 0.144288 0.000000 0.081315 0.000000 0.148014 0.000000 0.359863 0.000000
@@
Dynamic Linear ETH N/A 0.151146 0.000000 0.113141 0.000000 0.136730 0.000000 0.357417 0.000000
Static Linear ETH N/A 0.164609 0.000000 0.120294 0.000000 0.114837 0.000000 1.300485 0.000000
@@
Dynamic Linear BNB N/A 0.148749 0.000000 0.100731 0.000000 0.076976 0.000000 0.331775 0.000000
Static Linear BNB N/A 0.159099 0.000000 0.108706 0.000000 0.064113 0.000000 0.329707 0.000000
@@
Dynamic Linear ADA N/A 0.203062 0.000000 0.240034 0.000000 0.110527 0.000000 0.327942 0.000000
Static Linear ADA N/A 0.213023 0.000000 0.246243 0.000000 0.088722 0.000000 0.403005 0.000000
@@
Dynamic Linear XRP N/A 0.200807 0.000000 0.274879 0.000000 0.095273 0.000000 0.352569 0.000000
Static Linear XRP N/A 0.211918 0.000000 0.293106 0.000000 0.056372 0.000000 0.361912 0.000000
@@
A similar pattern is observed for Ethereum (ETH). The naive method records an absolute loss of 0.1778 and a quadratic loss of 0.1162. The StaticVWAP model (with absolute loss optimization) reduces these figures to 0.1386 and 0.0765, respectively, while the quadratic loss optimization further lowers the quadratic loss to 0.0734. The dynamic linear and static linear regression models yield mixed results. For example, the dynamic linear regression on ETH achieves a slightly lower $R^{2}$ (0.1367) than the StaticVWAP model, but its absolute and quadratic losses remain higher. For BNB, ADA, and XRP, similar trends are observed. The StaticVWAP model generally produces lower absolute and quadratic VWAP losses compared to the naive benchmark and static linear regression. Notably, for ADA—a more volatile asset—the StaticVWAP model optimized with absolute loss delivers an absolute loss of 0.1876 and a quadratic loss of 0.1570, outperforming both the naive approach and the linear regression variants. Similarly, for XRP, the StaticVWAP model achieves an absolute loss of 0.1824 and a quadratic loss of 0.1875, which are improvements over the baseline values (0.2239 and 0.2759, respectively).
Regarding training times, the deep-learning models (StaticVWAP) generally require between 15 and 18 seconds per run, which, although longer than the near-instantaneous training times of linear regression models, represents an acceptable trade-off given the significant performance gains. Additionally, the dynamic approaches based on linear regression show mixed performance: although they sometimes yield positive $R^{2}$ values—suggesting a better fit to the volume curve—their VWAP losses are inconsistent across assets, with instances of complete failure on certain assets.
Overall, the results confirm that the proposed StaticVWAP model consistently outperforms the naive flat allocation and static linear regression benchmarks across multiple cryptocurrencies. Although the fixed volume curve approach achieves relatively good performance in matching the volume profile (as indicated by its $R^{2}$ scores), it does not translate into the lowest VWAP losses. The dynamic strategies provide some benefit over static linear regression; however, when applied to the deep-learning model, the dynamic framework does not improve performance and, in some cases, degrades it. These observations underscore that directly optimizing the VWAP execution objective—without relying solely on volume curve predictions—is crucial for achieving superior execution performance in volatile markets.
@@ 6.6 Training Times and Comparison Across Prediction Horizons
Tables 3 and 4 provide additional insight into the practical applicability of each method under different prediction horizons. Although the StaticVWAP models generally require longer training times than simpler models, their training durations remain computationally efficient—typically under 20 seconds—making the trade-off acceptable given the substantial improvements in VWAP loss performance. Notably, while the fixed volume curve approach converges rapidly on shorter horizons, its training time increases exponentially with longer prediction horizons.
When comparing the results for the 6-step-ahead allocation (Table 3), the relative ranking of the models remains consistent with the 12-step results, albeit with somewhat reduced absolute improvements. For instance, the deep-learning StaticVWAP model shows an improvement in absolute VWAP loss of approximately 13–16% relative to the naive benchmark, while the fixed volume curve approach achieves improvements in the range of 11–15%. In terms of quadratic VWAP loss, the deep-learning StaticVWAP model improves performance by roughly 38–40%, compared to improvements of 35–41% achieved by the fixed volume curve approach. In contrast, both static and dynamic linear regression models tend to yield modest improvements in absolute VWAP loss and often increased loss when considering the quadratic deviation. For the extended 48-step-ahead allocation (Table 4), the performance differences become more pronounced. The deep-learning StaticVWAP model delivers absolute VWAP loss improvements of between 16% and 28% and quadratic loss improvements ranging from 39% to 50% relative to the naive approach. The fixed volume curve approach, while still competitive with improvements of 15–21% (absolute loss) and 32–47% (quadratic loss), generally does not reach the performance levels of the StaticVWAP model. Moreover, the static and dynamic linear regression models show higher variability across different assets; for the most liquid assets they are able to improve both the absolute and quadratic deviation, while for the less liquid asset they tend to increase the quadratic loss, despite sometimes achieving the highest $R^{2}$.
Several key observations emerge from these results. With shorter prediction horizons, the performance gap between the fixed volume curve and the deep-learning StaticVWAP approaches narrows, likely due to easier optimization convergence with fewer prediction steps. In these scenarios, the dynamic framework applied to linear regression often yields poorer results than the static approach, which aligns with the original design of the dynamic framework for longer horizons. In contrast, with extended prediction horizons, the deterioration in performance when applying the dynamic framework to the StaticVWAP model becomes more pronounced—often falling below that of even simple linear regression—while the dynamic framework appears to benefit linear regression to a greater extent.The robustness of the deep-learning approach is further underscored by the consistently low standard deviations of the performance metrics relative to their means, indicating stable convergence regardless of initialization parameters. Overall, these observations confirm that the proposed method offers consistent calibration capability across assets, lookback periods, and prediction horizons, highlighting the advantage of directly integrating the execution strategy as an optimization objective rather than relying solely on volume curve predictions.
@@ Appendix A Results Tables
Dynamic Linear BTC N/A 0.098210 0.000000 0.045269 0.000000 -0.040100 0.000000 0.034625 0.000000
Static Linear BTC N/A 0.101460 0.000000 0.047394 0.000000 0.035099 0.000000 0.032666 0.000000
@@
Dynamic Linear ETH N/A 0.112606 0.000000 0.064005 0.000000 0.006629 0.000000 0.025424 0.000000
Static Linear ETH N/A 0.117375 0.000000 0.067607 0.000000 0.026996 0.000000 0.033748 0.000000
@@
Dynamic Linear BNB N/A 0.105881 0.000000 0.060695 0.000000 0.034713 0.000000 0.027022 0.000000
Static Linear BNB N/A 0.109159 0.000000 0.059817 0.000000 0.050791 0.000000 0.024749 0.000000
@@
Dynamic Linear ADA N/A 0.143405 0.000000 0.121989 0.000000 -0.000604 0.000000 0.028840 0.000000
Static Linear ADA N/A 0.147375 0.000000 0.123984 0.000000 0.024974 0.000000 0.028879 0.000000
@@
Dynamic Linear XRP N/A 0.134100 0.000000 0.139278 0.000000 0.062711 0.000000 0.029293 0.000000
Static Linear XRP N/A 0.138093 0.000000 0.144271 0.000000 0.043018 0.000000 0.032220 0.000000
@@
Dynamic Linear BTC N/A 0.237939 0.000000 0.186814 0.000000 0.254591 0.000000 1.139905 0.000000
Static Linear BTC N/A 0.259229 0.000000 0.197585 0.000000 0.186233 0.000000 0.678431 0.000000
@@
Dynamic Linear ETH N/A 0.249949 0.000000 0.262389 0.000000 0.215072 0.000000 1.561179 0.000000
Static Linear ETH N/A 0.279120 0.000000 0.282128 0.000000 0.153775 0.000000 0.999470 0.000000
@@
Dynamic Linear BNB N/A 0.245283 0.000000 0.203534 0.000000 0.136387 0.000000 0.992567 0.000000
Static Linear BNB N/A 0.269164 0.000000 0.230699 0.000000 0.097216 0.000000 0.860453 0.000000
@@
Dynamic Linear ADA N/A 0.366712 0.000000 0.692776 0.000000 0.150097 0.000000 1.235840 0.000000
Static Linear ADA N/A 0.394132 0.000000 0.704408 0.000000 0.096127 0.000000 1.080062 0.000000
@@
Dynamic Linear XRP N/A 0.358663 0.000000 0.571102 0.000000 0.130708 0.000000 1.857334 0.000000
Static Linear XRP N/A 0.387054 0.000000 0.623145 0.000000 0.059751 0.000000 0.527437 0.000000

### v2

@@ 6.3 Experimental Setup
The strategies are evaluated using hourly data from five major cryptocurrencies (BTC, ETH, BNB, ADA, and XRP) traded on Binance perpetual contracts covering the period from January 1, 2020, to July 1, 2024. This dataset provides sufficient historical data across varying market conditions to evaluate the models’ performance in both high and low volatility regimes. The data is partitioned chronologically, with the last 20% reserved as the test set to ensure a true out-of-sample evaluation. A validation set, comprising 20% of the remaining data, is randomly selected from the first 80% of the period to guide model selection and prevent overfitting to any specific market regime.
For input features, several categories are incorporated: (1) Volume data, which is transformed to address non-stationarity by dividing each volume value by a trailing two-week moving average, calculated with appropriate shifting to avoid look-ahead bias; (2) Temporal indicators including hour of day and day of week as categorical features to capture seasonality patterns; and (3) Price information in the form of returns calculated on the VWAP price of each bin, with a value of 0 assigned when no volume is traded.
The preparation of target variables differs based on the model type. For the neural network approaches, the normalization to volume curve is implemented directly within the loss function, with volumes divided by the sum over the lookback period for numerical stability. For traditional linear regression models, the volume curve is explicitly normalized by dividing each volume by the sum of volumes in the lookahead period, as these models cannot incorporate this normalization within their loss functions.
In these experiments, a lookback period of 120 time steps is used to predict allocations for 12 periods ahead. Additional experiments with modified parameters—one with a 6-step prediction horizon and a shorter lookback period, and another with a 48-step prediction horizon—are provided in the Appendix. The results from these experiments confirm that the framework is effective across different timeframes and that the underlying concepts generalize well.
@@ 6.5 Empirical Results
Table 2 summarizes the performance of various VWAP execution strategies for a 12-step-ahead allocation with a 120-step lookback window. The table reports the mean and standard deviation of the Absolute VWAP Loss (scaled by $10^{-2}$), Quadratic VWAP Loss (scaled by $10^{-4}$), the $R^{2}$ score for the volume curve, and the training time (in seconds) for each strategy, across five assets (BTC, ETH, BNB, ADA, and XRP).
Several key patterns emerge from these results. First, all proposed methods consistently outperform the naive flat allocation approach across all assets, confirming that more sophisticated allocation strategies yield tangible benefits for VWAP execution.
@@
Dynamic Linear BTC N/A 0.142535 0.000000 0.085731 0.000000 0.193757 0.000000 0.241681 0.000000
Static Linear BTC N/A 0.146979 0.000000 0.084170 0.000000 0.162770 0.000000 0.244588 0.000000
@@
Dynamic Linear ETH N/A 0.158829 0.000000 0.121279 0.000000 0.169994 0.000000 0.249899 0.000000
Static Linear ETH N/A 0.166326 0.000000 0.122063 0.000000 0.135824 0.000000 0.278128 0.000000
@@
Dynamic Linear BNB N/A 0.160003 0.000000 0.115190 0.000000 0.127731 0.000000 0.262799 0.000000
Static Linear BNB N/A 0.163049 0.000000 0.114673 0.000000 0.101505 0.000000 0.216979 0.000000
@@
Dynamic Linear ADA N/A 0.213691 0.000000 0.258357 0.000000 0.142651 0.000000 0.236307 0.000000
Static Linear ADA N/A 0.217345 0.000000 0.253104 0.000000 0.112113 0.000000 0.293910 0.000000
@@
Dynamic Linear XRP N/A 0.213728 0.000000 0.295177 0.000000 0.098195 0.000000 0.290283 0.000000
Static Linear XRP N/A 0.216581 0.000000 0.295605 0.000000 0.071538 0.000000 0.248654 0.000000
@@
When comparing the methods based on their ability to predict volume curves, we observe that Static Linear Regression is more effective than the StaticVWAP model when the latter is optimized for absolute or quadratic VWAP loss. However, the Dynamic Linear Regression approach consistently achieves the highest $R^{2}$ values across all assets, indicating superior volume curve prediction capability. This improved prediction translates into better VWAP execution compared to Static Linear Regression.
Interestingly, despite Dynamic Linear Regression’s advantage in volume prediction, the Fixed Volume Curve approach demonstrates superior VWAP execution performance. This finding underscores a critical insight: accurate volume curve prediction, while beneficial, is not the primary determinant of optimal VWAP execution.
Most significantly, the proposed StaticVWAP deep learning model consistently achieves the best absolute or quadratic VWAP loss values (depending on its optimization target) across all assets. When trained to minimize absolute VWAP loss, it delivers the lowest absolute slippage; similarly, when optimized for quadratic VWAP loss, it produces the lowest quadratic slippage metrics.
Regarding training times, the deep-learning models (StaticVWAP) generally require between 15 and 18 seconds per run, which, although longer than the near-instantaneous training times of linear regression models, represents an acceptable trade-off given the significant performance gains. The dynamic approaches show mixed results when applied to the StaticVWAP model, often degrading performance rather than improving it.
Overall, these results validate the central thesis of this study: directly optimizing the VWAP execution objective—rather than focusing solely on volume curve prediction—is essential for achieving superior execution performance in volatile markets such as cryptocurrencies. The StaticVWAP deep learning framework provides a flexible and effective approach that consistently outperforms conventional methods, even when utilizing a relatively simple model architecture.
@@ 6.6 Training Times and Comparison Across Prediction Horizons
Tables 3 and 4 provide additional insight into the practical applicability of each method under different prediction horizons. Several consistent patterns emerge across these timeframes that reinforce the findings from our main experiments.
First, regarding training efficiency, the StaticVWAP models consistently maintain reasonable computational demands across all horizons—typically under 20 seconds—making them practical for real-world applications despite being more complex than linear regression models. Notably, while the fixed volume curve approach converges rapidly for shorter horizons, its training time increases dramatically with longer prediction horizons, eventually requiring several minutes per run for 48-step predictions.
When examining performance across different timeframes, the relative rankings of the models remain remarkably consistent. For both shorter (6-step) and longer (48-step) prediction horizons, the StaticVWAP models optimized directly for VWAP loss consistently outperform both naive approaches and linear regression variants. The fixed volume curve approach remains competitive, particularly when optimized with quadratic loss, but generally does not reach the performance levels of the StaticVWAP model.
Dynamic approaches show consistent patterns across different horizons. Dynamic linear regression demonstrates increasingly strong volume curve prediction capability (higher $R^{2}$ values) as the prediction horizon extends, yet this improvement in volume prediction does not translate to proportional gains in VWAP execution performance. This pattern strongly supports our central thesis that accurate volume prediction, while informative, is not the primary driver of optimal VWAP execution. In fact, the performance gap between methods directly optimized for VWAP execution and those focused on volume prediction tends to persist or even widen with longer horizons.
The stability of these findings across different prediction horizons and assets underscores the robustness of our approach. Whether executing shorter or longer-term VWAP strategies, directly optimizing the execution objective consistently yields superior results compared to approaches that rely primarily on volume curve prediction.
@@ Appendix A Results Tables
Dynamic Linear BTC N/A 0.100858 0.000000 0.048049 0.000000 0.109054 0.000000 0.049620 0.000000
Static Linear BTC N/A 0.102500 0.000000 0.048115 0.000000 0.087037 0.000000 0.035565 0.000000
@@
Dynamic Linear ETH N/A 0.114201 0.000000 0.063881 0.000000 0.103432 0.000000 0.034144 0.000000
Static Linear ETH N/A 0.117400 0.000000 0.065136 0.000000 0.075838 0.000000 0.024283 0.000000
@@
Dynamic Linear BNB N/A 0.109450 0.000000 0.060628 0.000000 0.083574 0.000000 0.021749 0.000000
Static Linear BNB N/A 0.111120 0.000000 0.060123 0.000000 0.061241 0.000000 0.021126 0.000000
@@
Dynamic Linear ADA N/A 0.145723 0.000000 0.123286 0.000000 0.089187 0.000000 0.018582 0.000000
Static Linear ADA N/A 0.147620 0.000000 0.120900 0.000000 0.065587 0.000000 0.018901 0.000000
@@
Dynamic Linear XRP N/A 0.140321 0.000000 0.150192 0.000000 0.066619 0.000000 0.019342 0.000000
Static Linear XRP N/A 0.141379 0.000000 0.148847 0.000000 0.047692 0.000000 0.023341 0.000000
@@
Dynamic Linear BTC N/A 0.255156 0.000000 0.214156 0.000000 0.259255 0.000000 0.266352 0.000000
Static Linear BTC N/A 0.265879 0.000000 0.206610 0.000000 0.193961 0.000000 0.262161 0.000000
@@
Dynamic Linear ETH N/A 0.263310 0.000000 0.289951 0.000000 0.229208 0.000000 0.265109 0.000000
Static Linear ETH N/A 0.284703 0.000000 0.289589 0.000000 0.160990 0.000000 0.340382 0.000000
@@
Dynamic Linear BNB N/A 0.270936 0.000000 0.245040 0.000000 0.163863 0.000000 0.244099 0.000000
Static Linear BNB N/A 0.279849 0.000000 0.247151 0.000000 0.109303 0.000000 0.247065 0.000000
@@
Dynamic Linear ADA N/A 0.388671 0.000000 0.792817 0.000000 0.172699 0.000000 0.251263 0.000000
Static Linear ADA N/A 0.405453 0.000000 0.754683 0.000000 0.107097 0.000000 0.249000 0.000000
@@
Dynamic Linear XRP N/A 0.391767 0.000000 0.675384 0.000000 0.135234 0.000000 0.237720 0.000000
Static Linear XRP N/A 0.400026 0.000000 0.669650 0.000000 0.077930 0.000000 0.268407 0.000000